Alexis spent $8.28 for two bags of dog biscuits. If there are 36 biscuits per bag, find the unit cost per biscuit.
Found 2 solutions by ewatrrr, learningalgebra13:
Answer by ewatrrr(24785)


Hi,
2 bags, 36 biscuits/bag
2*36 = 72 biscuits in all
$8.28/72 = 11.5 cents/biscuit
CHECKING our Answer
$0.115*72 = $8.28

Answer by learningalgebra13(2)

Money
-------
Biscuits

8.28 x
---- = ----
2 72
cross multiply
2*x=2x
8.28*72=5.96
plug it in
2x=5.96
divid 2 to get x by its self
/2 /2
x=2.98